# Safe MPC Alignment with Human Directional Feedback

### v1

@@ V-D Detailed Algorithm of Safe MPC Alignment
With the detailed centering choice method of $\boldsymbol{\theta}_{i}$ as well as the main results on convergence and misspecification, we can establish the full safe MPC alignment. With initializing $\Theta_{0}$ and pre-define a termination threshold $\rho_{H}$, three steps in the hypothesis space cutting are executed until user satisfies robot motion or the iteration count reaches the maximum count $K$ in (25). Note that during the execution of the MPC policy, $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the human can perform emergency stopping and reset the robot’s state, which corresponds to Section III-C3.
When reaching and exist after $K-1$ iterations, the algorithm checks the misspecification by (26). The Safe MPC Alignment algorithm with misspecification identification is presented in the Algorithm 1.
@@
VI Numerical Example and User Study in Simulation Games
In this section, we will first evaluate the proposed Safe MPC Alignment method using a numerical example with simulated human corrections knowing ground-truth safety constraints. Then, we will test the performance of the methods in two user studies using two simulation games developed based on MuJoCo [61].We consider two typical robot tasks. One is a drone navigation task and the other a 7-DOF robot arm reaching task. In both user studies, the robot is controlled by an MPC controller and human participants provide simulation robots with real-time correction instructions through a keyboard. The objective of each game is to teach a robot to successfully learn a safe MPC to move through an unknown safety-critical environment.
The organization in this section is as follows: Subsection VI-A presents the numerical evaluation and its analysis. Subsections VI-B and VI-C describe the setting of the two developed simulation games used in our user study. Subsection VI-D details the study procedure and the recruitment of human users. The results of the user studies and some analysis are given in VI-E. Please access the accompanying codes at https://github.com/asu-iris/Safe-MPC-Alignment.
VI-A Numerical Example
VI-A1 Environment Setup
We use the safe control of an inverted pendulum as the numerical example. The dynamics equation of an inverted pendulum is
@@ VI-A1 Environment Setup
with $\alpha$ being the angle between the pendulum and the direction of gravity, $\mathrm{g}=10\mathrm{m/s^{2}}$ the gravity constant, $u$ the pivot torque, $m=1\mathrm{kg}$, $l=1\mathrm{m}$ and $d=0.4$ being the length, inertia, and dampening of the pendulum. We set the system state as $\boldsymbol{x}=[\alpha,\dot{\alpha}]^{\mkern-1.5mu\mathsf{T}}$ and discretize (27) by the Euler method with $\triangle t=0.02\mathrm{s}$. A zero-mean Gaussian noise with covariance $\Sigma=\mathbf{diag}(1,4)\times 10^{-5}$ is added to $\boldsymbol{x}$ at each control step to simulate the state noise. The task is to control the pendulum to the target $\boldsymbol{x}^{*}=[\pi,0]^{\mkern-1.5mu\mathsf{T}}$. We additionally pose the following unknown state affine constraint to the motion of the pendulum
@@
where $\boldsymbol{\theta}^{H}=[0.6,1]^{\mkern-1.5mu\mathsf{T}}$ is to be learned.
In Safe MPC formulation (1), we set $T=40$, the step cost $c(\boldsymbol{x}_{t},u_{t})$ and terminal cost $h(\boldsymbol{x}_{T})$
$\begin{gathered}c(\boldsymbol{x}_{t},u_{t})=Qu_{t}^{2}\\ h(\boldsymbol{x}_{T})=(\boldsymbol{x}_{T}-\boldsymbol{x}^{*})^{T}R(\boldsymbol{x}_{T}-\boldsymbol{x}^{*}),\end{gathered}$ (29)
respectively, with $Q=0.1$ and $R=\mathrm{diag}(25,10)$.
To ensure the safe MPC rollout trajectory to satisfy a global constraint, it is sufficient to learn a safety constraint in safe MPC defined on the closest controllable state. Thus, in (1), we aim to define and learn a constraint defined on the first controlled state:
@@
In asymptotic MPC approximation, we set $\gamma=0.1$.
Suppose the user intent set $\bar{\Theta}_{H}=\{\boldsymbol{\theta}|\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02\}$. Then, simulated human corrections $\mathbf{a}$ can be generated as follows: when the system is running near the human safety boundary, i.e., ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ ($\epsilon_{g}=0.25$ here), a correction $\mathbf{a}$ can be generated in probability $p=0.3$ using the first element the signed gradient vector $\mathrm{sign}(-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H}))$. Recall that $B(\boldsymbol{\xi},\boldsymbol{\theta})$ is the barrier function in (2). It is trivial to verify that this correction $\mathbf{a}$ has positive product with $-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$, and by definition it is applied to a safe trajectory $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$. The both hypotheses in section IV-A is then satisfied, and $\mathbf{a}$ can be used as a valid human correction.
@@ VI-A2 Convergence Analysis
In the first session, we evaluate the convergence of the learning algorithm without hypothesis misspecification. In Algorithm 1, we set the initial hypothesis $\Theta_{0}$ with box boundaries $\boldsymbol{c}_{l}=[-6,-6]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{h}=[2,2]^{\mkern-1.5mu\mathsf{T}}$; thus $\Theta_{0}$ includes $\bar{\Theta}_{H}$. The initial value of $\boldsymbol{\theta}_{1}$ is set as $[-2,-2]^{\mkern-1.5mu\mathsf{T}}$ (i.e., the center of $\Theta_{0}$).
With simulated human corrections, we run Algorithm 1 and track the update of $\boldsymbol{\theta}_{i}$ after every correction. $\boldsymbol{\theta}_{i}$ falls in the intention set $\bar{\Theta}_{H}$ after 14 corrections, triggering the termination of learning. We then respectively evaluate each history $\boldsymbol{\theta}_{i}$ by using it as the MPC constraint parameter to control the system starting from $\boldsymbol{x}=[0,0]^{\mkern-1.5mu\mathsf{T}}$. We plot the corresponding MPC rollout trajectory in $(\alpha,\dot{\alpha})$ phase plot in Fig. 8.The affine ground-truth constraint (28) and the corresponding learned constraints $\boldsymbol{g}_{\boldsymbol{\theta}_{i}}$ are also shown in Fig. 8.
@@
From Fig. 8, we can conclude that the ground-truth constraint can be successfully recovered from the simulated corrections. The precision of this recovery is very high and the learned and ground-truth safety boundary almost overlap.
To further analyze the convergence of the algorithm, we run the algorithm for 10 times, all of which terminate after 12-17 corrections. In each run, we keep track of the following two metrics over correction count, shown in Fig. 9.
M1 (left panel of Fig. 9): The logarithm volume measure of the solved MVE $\log\det\bar{H}$ in (21). This provides a lower bound of $\Theta_{i}$ to show the exponential convergence property.
M2 (right panel of Fig. 9): The euclidean distance between learned parameter $\boldsymbol{\theta}_{i}$ and $\boldsymbol{\theta}_{H}$, i.e $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$. This suggests the similarity between the learned constraint and the ground-truth constraint, see Fig. 9.
From Fig. 9, we can conclude the proposed algorithm is
@@
    Fast-converging. During learning, the logarithm volume $\log\det\bar{H}$ has near-linear decreasing property. This observation is consistent with the exponential convergence in theory stated in Lemma 5.
@@
    Correction-efficient. During learning, $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ converges to the value lower than 0.02 within 17 corrections. By using 2D area as a volume measure, we can compute the upper bound of correction count by Theorem 1 :
@@
    This suggests that the algorithm is guaranteed to converge within finite $K=17$ corrections, this is consistent with the observations in the experiment.
@@
(b) Loss to Ground-truth
@@
VI-A3 Result Analysis for Misspecification Case
To analyze the performance of the proposed algorithm in misspecification cases, we set the initial hypothesis $\Theta_{0}$ as $\hat{\boldsymbol{c}_{l}}=[-1,-1]^{\mkern-1.5mu\mathsf{T}}$ and $\hat{\boldsymbol{c}_{h}}=[0.8,0.8]^{\mkern-1.5mu\mathsf{T}}$ which does not contain $\bar{\Theta}_{H}$. In this setting, the algorithm automatically stops after 7th correction. The path of $\boldsymbol{\theta}_{i},\ i=1,\dots,8$ is shown in Fig. 10. In Fig. 10, $\boldsymbol{\theta}_{i}$ is continuously moving towards the boundary of the initial hypothesis space $\Theta_{0}$. This fact verifies Theorem 2 and justifies the proposed method is capable of misspecification detection.
@@ VI-A3 Result Analysis for Misspecification Case
VI-B Drone Navigation Game
VI-B1 Environment Setup
As shown in Fig. 11, the goal of this drone navigation game is let a human user to teach a done to successfully navigate through a narrow gate (rectangle frame in yellow) whose position is unknown, given different flying target positions (shown in yellow points behind the gate). The drone starts from $[x_{\text{init}},y_{\text{init}},z_{\text{init}}]^{\mkern-1.5mu\mathsf{T}}$ with $x_{\text{init}}=0.1$, $z_{\text{init}}=0.5$, and $y_{\text{init}}$ uniformly randomly in an interval $[4,6]$. The starting quaternion $\boldsymbol{q}_{\text{init}}=[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$. We set three different target positions at $[19,1,9]^{\mkern-1.5mu\mathsf{T}}$, $[19,5,9]^{\mkern-1.5mu\mathsf{T}}$ and $[19,9,9]^{\mkern-1.5mu\mathsf{T}}$, with target quaternion $[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$. For the narrow gate (yellow rectangle), both its size and position are unknown, and one only roughly know its x position $x_{\text{gate}}\approx 10.0$. The drone navigation game is developed using MuJoCo physics engine [61].

### v2

@@ V-D Detailed Algorithm of Safe MPC Alignment
With the main results on convergence and misspecification, we can establish the full safe MPC alignment algorithm in Algorithm 1. With initializing $\Theta_{0}$ and pre-define a termination threshold $\rho_{H}$, three steps in the hypothesis space cutting are executed until the user satisfies robot motion or the iteration count reaches the maximum count $K$ in (24). Note that during the execution of the MPC policy, $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the human can perform emergency stopping and reset the robot’s state, which corresponds to Section III-C3. After $K-1$ iterations, the algorithm checks the misspecification by (25).
@@
VI Numerical Examples and Analysis
In this section, we will evaluate the proposed Safe MPC Alignment using two numerical examples: inverted pendulum and two-link torque-controlled robot arm moving in constrained environment. Simulated human corrections knowing ground-truth safety constraints will be used. We will compare the proposed method with two baselines, and conduct ablation studies on the hyperparameters of the algorithm. All numerical experiments are implemented in Python, running on a Dell desktop with Intel i9-13900K CPU. All our experiment codes are open-sourced at at https://github.com/asu-iris/Safe-MPC-Alignment.
VI-A Inverted Pendulum
VI-A1 Setup
The dynamics of an inverted pendulum is
@@ VI-A1 Setup
with $\alpha$ being the angle between the pendulum and gravity, $\mathrm{g}$ the gravity constant, $u$ input torque, $m$, $l$ and $d$ the length, inertia, and damping ratio. We set the system state as $\boldsymbol{x}=[\alpha,\dot{\alpha}]^{\mkern-1.5mu\mathsf{T}}$ and discretize (26) by the Euler method with $\triangle t=0.02\mathrm{s}$. A zero-mean Gaussian noise with covariance $\Sigma=\mathbf{diag}(1,4)\times 10^{-5}$ is added to $\boldsymbol{x}$ at each control step to simulate the state noise. The task is to control the pendulum to the target $\boldsymbol{x}^{*}=[\pi,0]^{\mkern-1.5mu\mathsf{T}}$. We additionally pose an unknown global state constraint on pendulum motion:
@@
where $\boldsymbol{\theta}^{H}=[0.6,1]^{\mkern-1.5mu\mathsf{T}}$ is to be learned. All other parameter values are given in Appendix.
In Safe MPC (1), we set $T=40$ and the cost functions as
$\begin{gathered}c(\boldsymbol{x}_{t},u_{t})=Ru_{t}^{2}\\ h(\boldsymbol{x}_{T})=(\boldsymbol{x}_{T}-\boldsymbol{x}^{*})^{T}Q(\boldsymbol{x}_{T}-\boldsymbol{x}^{*}),\end{gathered}$ (28)
respectively, with $R=0.1$ and $Q=\mathrm{diag}(25,10)$. To ensure the safe MPC policy to satisfy a global constraint, it is sufficient to learn a safety constraint defined on the closest controllable state. Thus, in (1), we define
@@
In the penalty-based MPC (3), we set $\gamma=0.1$.
Suppose the user intent set $\bar{\Theta}_{H}=\{\boldsymbol{\theta}|\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02\}$. The simulated human corrections $\mathbf{a}$ are generated as follows: when the system state is near the safety boundary, i.e., ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ ($\epsilon_{g}=0.25$ here), a directional correction $\mathbf{a}$ is be generated in probability $p=0.3$ using $\mathrm{sign}(-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H}))$ (from which only the first entry is taken considering (6)). The correction is probabilistic to simulate the randomness of human providing corrections. It is easy to verify that the simulated $\mathbf{a}$ is valid in term of satisfying the hypotheses in section IV-A.
@@ VI-A2 Convergence Analysis
We first evaluate the convergence of our algorithm without hypothesis misspecification. In Algorithm 1, we set the initial hypothesis $\Theta_{0}$ with box boundaries $\boldsymbol{c}_{l}=[-6,-6]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[2,2]^{\mkern-1.5mu\mathsf{T}}$; thus $\Theta_{0}$ includes $\bar{\Theta}_{H}$. The initial $\boldsymbol{\theta}_{1}=[-2,-2]^{\mkern-1.5mu\mathsf{T}}$ (i.e., the center of $\Theta_{0}$).
With each simulated human correction, we track $\boldsymbol{\theta}_{i}$ in each learning update. Each update of ${\Theta}_{i}$ takes less than 30ms. $\boldsymbol{\theta}_{i}$ falls in the ground-truth set $\bar{\Theta}_{H}$ after 14 corrections. The MPC rollout trajectory (phase plot) with each $\boldsymbol{\theta}_{i}$ is shown in Fig. 8. The ground-truth constraint (27) and the learned constraints $\boldsymbol{g}_{\boldsymbol{\theta}_{i}}$ are also plotted in Fig. 8. From Fig. 8, we can conclude that the ground-truth constraint can be successfully learned from simulated corrections.
@@
To further analyze the learning convergence, we run the algorithm for 10 times. In each run, we keep track of two metrics over correction (iteration) count and show them in Fig. 9. The first metric (left panel) is the logarithm volume measure of the solved MVE $\log\det\bar{H}$ in (20); this provides a lower bound of $\Theta_{i}$ to show the exponential convergence property. The second (right panel) is the parameter error $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$. From Fig. 9, we can conclude the proposed method is
@@
    Fast-converging. The logarithm volume $\log\det\bar{H}$ has near-linear decreasing speed. This is consistent with the exponential convergence in theory stated in Lemma 5.
@@
    Correction-efficient. $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ converges to the value lower than 0.02 within 17 corrections. This is consistent with the theoretical upper bound of correction count from given by Theorem 1 :
@@
(b) Parameter Estimation Error
@@
VI-A3 Analysis for Misspecification Case
To analyze the performance of our algorithm to certify misspecification cases, we set the initial hypothesis space $\Theta_{0}$ as ${\boldsymbol{c}_{l}}=[-1,-1]^{\mkern-1.5mu\mathsf{T}}$ and ${\boldsymbol{c}_{u}}=[0.8,0.8]^{\mkern-1.5mu\mathsf{T}}$ which does not contain $\bar{\Theta}_{H}$. In this setting, the algorithm automatically stops after 7th correction. The path of $\boldsymbol{\theta}_{i},\ i=1,\dots,8$ is shown in Fig. 10. In Fig. 10, $\boldsymbol{\theta}_{i}$ is continuously moving towards the boundary of $\Theta_{0}$. This fact verifies Theorem 2 and justifies the proposed method is capable of misspecification detection.
@@ VI-A3 Analysis for Misspecification Case
VI-B Two-link Robot Arm Reaching with Complex Constraint
VI-B1 Setup
The dynamics if torque controlled two-link robot arm moving in the horizontal plane is $M(\boldsymbol{q})\ddot{\boldsymbol{q}}+\boldsymbol{c}(\boldsymbol{q},\dot{\boldsymbol{q}})=\boldsymbol{\tau}$, where $M(\boldsymbol{q})$ is the inertia matrix; $\boldsymbol{c}(\boldsymbol{q},\dot{\boldsymbol{q}})$ is the Coriolis term. $\boldsymbol{q}=[q_{1},q_{2}]^{\mkern-1.5mu\mathsf{T}}$ defines the two joint angles of the robot. The system state is defined as $\boldsymbol{x}=[\boldsymbol{q},\dot{\boldsymbol{q}}]$. The control input $\boldsymbol{u}$ is the joint torque $\boldsymbol{\tau}=[\tau_{1},\tau_{2}]^{\mkern-1.5mu\mathsf{T}}$, i.e. $\boldsymbol{u}=\boldsymbol{\tau}$. The other detailed physical parameters are given in Appendix.
As shown in Fig. 12, the robot starts at $\boldsymbol{x}_{\mathrm{init}}=[-\frac{\pi}{2},0,0,0]^{\mkern-1.5mu\mathsf{T}}$ and the task is to reach a target $\boldsymbol{x}^{*}=[\frac{\pi}{2},0,0,0]^{\mkern-1.5mu\mathsf{T}}$. In Safe MPC (1), we set $T=30$, the step cost functions
$\begin{gathered}c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})=\boldsymbol{x}_{t}^{\mkern-1.5mu\mathsf{T}}P\boldsymbol{x}_{t}+\boldsymbol{u}_{t}^{\mkern-1.5mu\mathsf{T}}R\boldsymbol{u}_{t},\\ h(\boldsymbol{x}_{T})=(\boldsymbol{x}_{T}-\boldsymbol{x}^{*})^{T}Q(\boldsymbol{x}_{T}-\boldsymbol{x}^{*}),\end{gathered}$ (31)
with $P=\mathrm{diag}(1,1,2,2)$, $R=\mathrm{diag}(0.1,0.1)$ and $Q=\mathrm{diag}(5,5,10,10)$.
As shown by the black lines (obstacles) in Fig. 12, we define a complex reaching environment with non-convex obstacles and narrow corridor. The ground-truth global safety constraint corresponding to Fig. 12 is given
${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi})=-1+\boldsymbol{\theta}_{H}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ (32)
with $\boldsymbol{\phi}{=}[\mathrm{tanh}(x_{e}y_{e}),\mathrm{tanh}(x_{e}^{2},\mathrm{tanh}(y_{e}^{2}),\mathrm{tanh}(x_{e}),\mathrm{tanh}(y_{e})]^{\mkern-1.5mu\mathsf{T}}$, $\boldsymbol{\theta}_{H}=[0.78,0.91,-0.52,0.52,-1.04]^{\mkern-1.5mu\mathsf{T}}$, and $[x_{e},y_{e}]^{\mkern-1.5mu\mathsf{T}}$ is the end-effector position of the robot arm. The human intent set is $\bar{\Theta}_{H}{=}\{\boldsymbol{\theta}|\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.2\}$. We set the initial hypothesis $\Theta_{0}$ with box boundaries $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[3,\dots,3]^{\mkern-1.5mu\mathsf{T}}$. The simulated human corrections are generated in the similar way to that in Section VI-A.
Figure 11: The heatmap of a learned constraint $g$. The learned and ground-truth constraints are almost identical, showing the success of our method for learning complex constraint from corrections.
VI-B2 Result
Our method can successfully learn the safety constraint within 25-30 simulated corrections. One instance of the learned constraint is shown in Fig. 11. The comparison of the robot reaching motion without and with the learned safety constraint is shown in Fig. 12. The robot is capable of operating in the intended safe region after the algorithm converges. Those results show that the efficacy of the proposed method for solving system with complex safety constraints.
(a) Trajectory Before Corrections
(b) Trajectory After Corrections
Figure 12: The trajectory of the two-link robot arm (a) without and (b) with learned safety constraint. The robot successfully achieves the task with the learned safety constraint.
VI-C Baseline Comparison
In this session, we compare our method with two baseline methods using magnitude corrections in the pendulum environment (the task and model settings are identical to Section VI-A). As reviewed previously, the direct literature learning safety constraints from human feedback is sparse to the authors’ best knowledge. Thus, the chosen baselines are an adaptation from the most relevant methods.
Baseline 1: gradient matching. This baseline is inspired by [59], where raw human correction (with magnitude) is directly used as the gradient direction. In a similar spirit, when a human feedback $\mathbf{a}$ is made on robot trajectory $\xi_{\boldsymbol{\theta}}^{\gamma}$, we directly use the human magnitude correction as the true gradient label to learn the constraint $\boldsymbol{\theta}$. This leads to a learning loss
$\mathcal{L}=\left\lVert\mathbf{a}+\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\right\rVert^{2}$ (33)
Baseline 2: maximum likelihood learning. This baseline follows [60], where it views $e^{{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})}$ as a probability of human applying the "stop-feedback" to robot motion $\boldsymbol{\xi}$, when the robot reaches in a $\epsilon_{g}$-proximity of the safety boundary, i.e., when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ Thus, a dataset of trajectory-correction paris, i.e., $\mathcal{D}=\{\mathbf{a}_{i},\boldsymbol{\xi}_{i}\}_{i=0}^{N}$, can be used as the training data to learn $\boldsymbol{\theta}$ using maximum likelihood, while also penalizing positivity of ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})$ as a regulizer term. This leads to the following loss
$\mathcal{L}=\sum\nolimits_{i=1}^{N}\big{(}-{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})+\alpha\cdot\mathrm{softplus}(g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i}))\big{)}$ (34)
In our comparison, we choose $N=15$ (similar to correction number needed for our method to converge) and $\alpha=5$ for its best performance.
The comparison between our method and Baseline 1 focuses on data efficiency. The convergence criteria of $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02$ is set for both methods. 50 independent runs are performed and the number of correction needed for convergence are counted. We use Adam optimizer [61] with learning rate 0.02. We summarize the results in Table I. The parameter estimation error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration (correction count) for a best run (i.e., best performance for the baseline) is plotted in Fig. 13(a). The comparison results demonstrate the significant data efficiency of the proposed method compared to gradient matching baseline, even though the latter uses both magnitude and direction information of corrections to learn.
TABLE I: Comparison of Safe MPC Alignment and Baseline I
Method Correction count for convergence
Safe MPC Alignment $11.76\pm 2.14$
Gradient Matching $240.38\pm 120.70$
(a) Comparison with Baseline 1
(b) Comparison with Baseline 2
Figure 13: (a) Comparison with gradient matching baseline: the parameter estimation error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration (correction count). This is best run (among 50 runs) with the least corrections for the baseline. Our method converges after 12 corrections, while the baseline needs more than 160 corrections to converge. (b) Comparison with the maximum likelihood baseline under the different $\epsilon_{g}$. In the plot, the y-axis is $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ at convergence and the error bar denotes the standard deviation. The maximum likelihood baseline fails to learn the correct weights when $\epsilon_{g}$ is large, while our method consistently shows high learning accuracy.
The comparison between our method and maximum likelihood baseline focuses on the varying positions of simulated correction, i.e., varying $\epsilon_{g}$. Recall that the simulated corrections will be generated only when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$. The value of $\epsilon_{g}$ determines the position of the correction applied. In our comparison, we vary $\epsilon_{g}$ from 0.1 to 0.9 and track the converged parameter error for both methods. As shown in Fig. 13(b), safe MPC alignment consistently converges to the ground-truth $\boldsymbol{\theta}_{H}$ regardless of $\epsilon_{g}$, while the maximum likelihood baseline fails when $|\epsilon_{g}|$ is large. The results suggest the robustness of the proposed method against different correcting conditions compared to maximum likelihood baseline method.
VI-D Ablation Study
In this session, we conduct two ablation studies to show the performance of our algorithm given different design parameters/aspects. The task is still based on pendulum environment with settings identical to Section VI-A).
VI-D1 Choosing $\boldsymbol{\theta}_{i}$: MVE centering versus random sampling
In the first study, we compare two strategies in selecting $\boldsymbol{\theta}_{i}$ from the hypothesis space $\Theta_{i-1}$ in our algorithm: MVE centering (used in our method) and random sampling. For the latter, $\boldsymbol{\theta}_{i}$ is chosen by sampling from the uniform distribution supported by $\Theta_{i-1}$. We compare the total number of corrections needed at learning convergence under the two strategies.
With 50 independent runs, we observe that the MVE centering requires $11.76\pm 2.14$ corrections, while random sampling requires $11.88\pm 3.50$ corrections. While the mean value of the correction count looks similar, the random sampling shows a slightly larger variance. This is understandable given the sampling randomness. We also observed that sampling process introduced large computational overhead, which is around 25% time slower than MVE centering. In contrast, the MVE centering used in our algorithm shows low variance and computational efficiency in finding $\boldsymbol{\theta}_{i}$. Because selecting $\boldsymbol{\theta}_{i}$ is casted as a convex programming for MVE centering, such advantage would be more significant for high-dimensional parameter space.
VI-D2 Choosing $\gamma$
In the second study, we investigate how the setting of $\gamma$ affects the learning performance. We vary $\gamma$ from 0.01 to 10. Fig. 14 shows the total number of corrections needed at convergence versus different $\gamma$s.
The result in Fig. 14 shows that the learning performance remains the same given different settings of $\gamma$. Since $\gamma$ is weights parameter to control the safety conservativeness in the penalty-based MPC policy in (3), it means our method can successfully learn a safety constraint when the robot executes with safe MPC policy of different conservativeness levels.
Figure 14: The total number of corrections needed for convergence under different $\gamma$, the x-axis is of log scale. Safe MPC alignment can converge with a similar number of corrections (iterations) under different levels of $\gamma$.
VII User Study in Simulation Games
In this section, we test the Safe MPC Alignment method in two real-human user studies using two simulation games developed based on MuJoCo [62]. One is the drone navigation game and the other the Franka robot arm reaching game. In both games, the robot is controlled by an MPC controller and human participants provide real-time corrections through a keyboard. The objective of each game is to teach a robot to successfully learn a safe MPC to move through environments with unknown obstacles. In following, Subsection VII-A describes the basic setting of the two games. Subsection VII-B details the user study procedure and the recruitment of human participants. The results and analysis are given in VII-C.
VII-A Setting of Two Simulation Games
In both user studies, human participants are asked to give correction to the following two physics-based simulation games developed based on MuJoCo.
(1) Drone Navigation Game. As shown in Fig. 15, the goal of this game is to let a drone to successfully navigate through a narrow gate (yellow rectangle) whose position is unknown, reaching different targets (shown in yellow points behind the gate). The safety constraint $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ is parameterized using a linear combination of spacial radial basis functions (RBFs), defined on y-z plane at $x=10$. The drone’s control inputs are the thrusts of four rotors and human correction is applied via keyboard, with specific keys mapping to specific correction commands to the drone’s control space. See the detailed task setup and keyboard interface in Appendix B-C.
(2) Franka arm reaching game. As in Fig. 16, the goal of the Franka arm reaching game is to teach a Franka arm such that its gripper hand successfully reaches different target positions (cyan ball) through an unknown narrow slot (two yellow bars). Note that the safety constraint is only applied on the end effector and the collision between the link and the slot is ignored. The safety constraint function $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ is parameterized as a weighted combination of RBFs defined on the $z$ axis and rotations around the $x$ axis. Human corrections are applied via keyboard, with specific keys mapping to specific correction commands applied to the robot control space $\mathbb{R}^{6}$ (we use operational space control). See the detailed task setup and keyboard interface in Appendix B-D.
In both games, the robots are controlled with MPC controller (3) but fail to maintain safety with the initial safety constraints. To simulate an emergency stop, the “Enter” key is used as a reset command, allowing users to reset the environment to its initial state whenever the participant deems the robot violating the safety constraint.